OpenStudy (anonymous):

Both triangles are similar. The area of the smaller triangle is about 270 ft^2. What is the best approximation for the larger triangle?

OpenStudy (anonymous):



OpenStudy (anonymous):

Ever seen this formula? A1 / A2 = (l1 / l2)^2

OpenStudy (anonymous):

Dont think i have

OpenStudy (anonymous):

Well, now you've seen it.

OpenStudy (anonymous):

Area 1 / area 2 = (length 1 / length 2)^2

OpenStudy (anonymous):

(25/35)^2= .510

OpenStudy (anonymous):

So A1 = 270, and 270 / A2 = .510

OpenStudy (anonymous):

Solve for A2 and thats your answer